Question 9 Find 5x+1 f(x) dx given that f(x) = { 3-x 0<x< 1 1<x<3 =la ole ala ala T
Answer #1

13-1 143 g emdr = J fan8x + J, 8ode - J5x+) + t- = 51)(x) + 01-() = [ J - 이 + [1-0] +33-D [ - ] =돌 146 올일 는 5-4월 +7 NE NW